# The role of porosity in the transition to inertial regime in porous media flows

### v2

@@ 3.2 Splitting the volume-integrated tortuosity $T_{\Omega}$
As noted previously, the proper identification of the percolating and recirculating volumes in the fluid is very sensitive to stray streamlines (i.e., those that, due to numerical errors, cross the boundary between the percolating and recirculating volumes). We observed that even several stray streamlines can significantly alter the value of the streamline-based tortuosity. Because of this, we conduct the numerical investigation of the two contributions to $T_{\Omega}$ on simple cubic (SC) systems, and use the obtained results to explain the behavior of $T_{\Omega}$ in the stochastic samples. SC samples can be thought of as models of particular fragments of a stochastic PM, at the same time having the geometry simple enough to allow for the accurate calculation of streamlines and identification of $\Omega_{p}$ and $\Omega_{v}$. To include in our analysis the cases with extremely high and low geometrical confinement, we consider SC systems of porosities $\phi=0.1$ and $\phi=0.999$. Exemplary visualizations of the steady-state velocity fields in the studied SC systems are shown in Figs. 14 and 15.
@@
Figure 14: Visualizations of the sections of the velocity field in $\phi=0.1$ SC system for $Re_{k}=2.61\cdot 10^{-7}$ (left) and $Re_{k}=0.049$ (right). The gray surfaces are the obstacles’ walls, and only a few chosen walls are shown for clarity.
@@
Figure 15: Visualizations of the sections of the velocity field in $\phi=0.999$ SC system for $Re_{k}=376$ (left) and $Re_{k}=1383$ (right). The gray surfaces are the obstacles’ walls.
3.3 Tortuosity as the inertia onset indicator in simple cubic systems
@@
We continue our analysis by looking at the Reynolds number dependence of $T_{\Omega}$ and $T_{s}$ for the SC sample, shown in Fig. 16. In the Darcy regime ($Re_{k}\rightarrow 0$), the values of both tortuosities stay constant and are compliant with each other, as shown in [27]. We attribute small visible discrepancies between the two tortuosities to the errors introduced during the generation of streamlines and the identification of percolating/recirculation volumes. When the Reynolds number increases, the streamline-based tortuosity, $T_{s}$, decreases for both systems, which was observed in stochastic and periodic porous media by other authors [22, 15]. On the other hand, when the volume-based tortuosity, $T_{\Omega}$, is concerned, the inertial effects manifest themselves differently depending on the porosity, similarly to the stochastic porous media (cf. Fig. 11). In the lower-porosity case, its value decreases slightly in the range $Re_{k}\in[5\cdot 10^{-4};3\cdot 10^{-3}]$ to later increase significantly until the steady-state regime ends. Such behavior resembles the one observed for the lower-porosity stochastic porous media in Fig. 11 and also for a real-life porous sample [16]. For the higher-porosity sample, the volume-based tortuosity follows the streamline-based one closely and diverges from it only at $Re_{k}\approx 10^{3}$. Similarly to the lower-porosity system, the initial decline of $T_{s}$ is followed by its rise, to a much smaller extent, however. This is in line with the behavior of $T_{\Omega}$ for the higher-porosity stochastic porous media shown in Fig. 11, and in [14].
@@
The strictly decreasing behavior of the streamline-based tortuosity can be attributed to the straightening of the percolating streamlines, which follows the growth of the recirculation zones volumes in the cavities and obstacles wakes. The volume-integrated tortuosity exhibits a more complex behavior, dependent on the sample’s porosity. Eq. (14) shows that the difference between $T_{\Omega}$ and $T_{s}$ is related to the integral of velocity magnitude inside the recirculation zones. Thus, we will further investigate the fall and the rise of the volume-based tortuosity in terms of the volume of the recirculation zones and the field of the velocity magnitude therein.
@@
Figure 16: $T_{\Omega}$ and $T_{s}$ for $\phi=0.1$ and $\phi=0.999$ SC samples. The inset in the left subplot shows the magnification of the $\phi=0.1$ plot to the range of $T_{s}$ values. The horizontal dashed lines represent the values of $T_{\Omega}$ calculated for vanishingly small $Re_{k}$.
@@
3.4 Origin of the $T_{\Omega}$ decrease and rise – volume of the recirculation zones
@@
Let us first focus on the numerator of the recirculation volume-related term from Eq. (14), $I[V]_{v}/I[V_{0}]_{p}$. For convenience, we divide it by the total fluid volume, $I[1]$, i.e.
@@
where the first term on the right-hand side represents the vortex volume normalized by the total volume of fluid, and the other one – the total velocity magnitude per unit recirculation zone volume. We consider each of those two components separately. Figs. 17 and 18 show the visualization of the percolating volumes and recirculation zones for low- and high-porosity SC samples, respectively, for two different Reynolds numbers each. In both systems, as $Re_{k}$ rises, the recirculation zones grow in volume, which in the case of the low-porosity system is visible through the percolating volume shrinking to a straight jet connecting the inlet and the outlet. To quantify the changes in the size of the recirculation zones, we plot their volume, $I[1]_{v}$, normalized by the total volume of fluid, $I[1]$, in Fig. 19. In the low-porosity system, there already exist recirculation zones in the Darcy regime (in the channels perpendicular to the main flow, cf. left visualization in Fig. 14). The increase in their volume starts at $Re_{k}\approx 5\cdot 10^{-4}$, which corresponds to the start of $T_{\Omega}$ decrease in Fig. 16. The vortex volume then continues to grow until the end of the steady state regime, with significant slowdown for $Re_{k}>10^{-2}$, the value at which the slope of the $T_{\Omega}$ rise in Fig. 16 decreases. This reduction in the recirculation volume growth pace indicates the strong geometrical confinement for the vortex development. We highlight that at the highest $Re_{k}$, there already exist secondary vortices in the perpendicular channels (cf. right visualization in Fig. 14). The total increase in the volume of the recirculation zone is about 56% of the fluid volume, or – equivalently – $0.56\cdot 0.1/0.9\approx 0.062$ of the volume of the solid obstacles. On the other hand, in the high-porosity system, there is no recirculation in the Darcy regime (cf. left visualization in Fig. 15). The increase of its volume starts at $Re_{k}\approx 2\cdot 10^{2}$, which is about two decades later than the visible start of $T_{\Omega}$ decrease for this system in Fig. 16. This discrepancy can be related to the ‘formation of inertial cores’ [44] when the inertia already starts to change the velocity field, although no separation yet occurs. As the recirculation volume starts to grow, it does so with an increasing slope until the end of the steady-state regime (cf. right visualization in Fig. 15). The increase in the volume of the recirculation zone in this case is about $0.63$ of the solid obstacle volume. The growth of the recirculation zones in the case of high-porosity SC systems was highlighted by other authors [19] as the reason for lower drag compared to the flow over a single obstacle. Finally, we note that as the considered SC systems have a regular and simple geometry, phenomena like the non-strictly-monotonic growth of the volume of the recirculation zones [45] related to the instantaneous reorganization of the percolating flow paths are not visible here. Nevertheless, the start of the recirculation zones volume growth can still be related to the deviation from Darcy’s law at $Re\approx 10^{-3}$ (compare the data for $\phi=0.1$ SC system in Figs. 19 and 21). Interestingly, no high-wavenumber oscillations can be seen in the values of $T_{\Omega}(Re_{k})$ in the case of the stochastic samples (cf. Fig. 11), which suggests that the reorganization of the percolating flow paths might not impact the velocity integrals of the volume-integrated tortuosity.
@@
Figure 17: Visualizations of the percolating volumes in $\phi=0.1$ SC system for the same values of $Re_{k}$ as in Fig. 14. The gray surfaces are the walls of the obstacles, and only a few chosen walls are shown for clarity. The visualizations use the histograms from which the percolating volumes were found. We show only the percolating volumes for clarity.
@@
Figure 18: Visualizations of the recirculating volumes for $\phi=0.999$ SC system for the same values of $Re_{k}$ as in Fig. 15. The gray surfaces are the walls of the obstacles. The visualizations use the histograms from which the recirculation volumes were found. We show only the recirculation volumes for clarity.
Figure 19: Relative vortex volumes, $I[1]_{v}/I[1]$, for the two studied SC systems.
@@
3.5 Origin of the $T_{\Omega}$ decrease and rise – energy confinement in the recirculation zones
@@
The increase of the recirculation zone volume can explain the decrease of $T_{\Omega}$ because it directly implies the straightening of all the percolating streamlines and thus the decrease of $T_{s}$ in Eq. 14. However, it fails to address the rise of the volume-based tortuosity at the highest Reynolds numbers. This is why now we turn our attention to the integral of the velocity magnitude inside the vortex per vortex volume, $I[V]_{v}/I[1]_{v}$ (further in the text, we shall refer to this quantity as the specific velocity magnitude in vortex, for brevity). Its dependence on the Reynolds number is shown in Fig. 20 for both SC systems. We note that since there is no recirculation zone present in the higher-porosity SC system in the Darcy regime, we assume that the value of the specific velocity magnitude in the vortex is zero in those cases. For both SC systems, the specific velocity magnitude in vortices increases with Reynolds number. In the low-porosity sample, the rise proceeds from the lowest $Re_{k}$ with the first-order slope ($p=1$) and experiences a jump in the slope around $Re_{k}=3\cdot 10^{-3}$. It later attains the slope of the order approximately one again, in the range of $Re_{k}$ corresponding to those where the growth of $T_{\Omega}$ slows down (cf. Fig. 16). The initial first-order dependence is explained by the linearity of the Stokes equation and proportionality between the body force magnitude $g$ and the mean flow velocity $\langle\boldsymbol{V}\rangle$ in the Darcy regime (i.e., when $g$ is increased, the velocity field inside the vortex is only scaled proportionally by a constant). The subsequently higher slope indicates that a non-trivial mechanism involving the momentum transfer to the recirculation zone and the growth of its volume comes into play in the inertial regime. At the same time, an analogy between the Darcy regime and the inertial regime above $Re_{k}=10^{-2}$ can be drawn in the sense that both are characterized by the recirculation zone growing more in kinetic energy than volume. This further highlights the importance of the geometrical confinement in the emergence of inertial effects in low-porosity samples [16]. In the high-porosity system, the start of the rapid growth of the specific velocity magnitude in the vortex happens roughly at $Re_{k}$ when the recirculation zone volume starts to grow (cf. Fig. 19). It is interesting that even though the integral of the velocity magnitude over the whole fluid volume, $I[V]$, achieved at the highest Reynolds number is about $40$ times larger in the high-porosity system (results not shown in the plots), the specific velocity magnitude in vortex is about 2 times higher in low-porosity system ($\approx 140$ vs. $\approx 70$). This suggests that strong geometrical confinement leads to the accumulation of relatively large amounts of the kinetic energy in the recirculation zones, quantified also by other authors using the variance of the kinetic energy [13, 15, 14, 16].
In the end, in the low-porosity stochastic and SC systems, the greater kinetic energy confinement in the recirculation zones and the cessation of their volume growth at high Reynolds numbers lead to the more pronounced increase of $T_{\Omega}$, compared to the higher-porosity counterparts.
@@
Figure 20: Integral of the absolute value of velocity in the vortex per vortex volume for SC systems (for brevity called the specific velocity magnitude in vortex in the text, $I[V]_{v}/I[1]_{v}$). The inset in the left subplot shows the same data but on a log-log scale.
4 Conclusions

### v3

@@ III.2 Tortuosity as the inertia onset indicator in simple cubic systems
To deeper investigate the origin of such different behavior of the tortuosities observed in the stochastic samples, we make use of the split of the volume-integrated tortuosity, Eq. (13), and consider the terms $T_{s}$ and $I[V]_{v}/I[V_{0}]_{p}$ separately. As noted previously, the proper identification of the percolating and recirculating volumes in the fluid is very sensitive to stray streamlines (i.e., those that, due to numerical errors, cross the boundary between the percolating and recirculating volumes). We observed that even several stray streamlines can significantly alter the value of the streamline-based tortuosity. Because of this, we conduct the numerical investigation of the two contributions to $T_{\Omega}$ on simple cubic (SC) systems, and use the obtained results to explain the behavior of $T_{\Omega}$ in the stochastic samples. SC samples can be thought of as models of particular fragments of a stochastic porous medium, at the same time having the geometry simple enough to allow for the accurate calculation of streamlines and identification of $\Omega_{p}$ and $\Omega_{v}$. In our analysis, we consider SC systems of porosities $\phi=0.1,0.59,0.93,0.999$, which include the cases with extremely high and low geometrical confinement. Exemplary visualizations of the steady-state velocity fields in the studied SC systems are shown in Figs. 1013.
@@
Figure 10: Visualizations of the sections of the velocity field in $\phi=0.1$ SC system for $Re_{k}=8.34\cdot 10^{-7}$ (top) and $Re_{k}=0.155$ (bottom). The gray surfaces are the obstacles’ walls, and only a few chosen walls are shown for clarity.
@@
Figure 11: Visualizations of the sections of the velocity field in $\phi=0.59$ SC system for $Re_{k}=3.72\cdot 10^{-14}$ (top) and $Re_{k}=8.70$ (bottom). The gray surfaces are the obstacles’ walls and black lines are chosen streamlines close to the obstacle’s surface.
Figure 12: Visualizations of the sections of the velocity field in $\phi=0.93$ SC system for $Re_{k}=1.73\cdot 10^{-12}$ (top) and $Re_{k}=2.01$ (bottom). The gray surfaces are the obstacles’ walls.
Figure 13: Visualizations of the sections of the velocity field in $\phi=0.999$ SC system for $Re_{k}=376$ (top) and $Re_{k}=1557$ (bottom). The gray surfaces are the obstacles’ walls.
We first look at the Reynolds number dependence of the volume-integrated tortuosity $T_{\Omega}$ for all SC samples in Fig. 14. As the porosity increases, the value of the volume-integrated tortuosity at vanishingly small Reynolds number, $T_{\Omega}(Re_{k}\rightarrow 0)$, decreases. This is an expected result, since in this regime $T_{\Omega}\equiv T_{s}$, and a smaller spherical obstacle means fewer streamlines significantly deviate from the straight shape. Secondly, looking at the minima of the functions in the bottom plot, one sees that as the porosity decreases, the value by which $T_{\Omega}$ falls before reaching the minimum decreases as well, relative to $T_{\Omega}-1$ at $Re_{k}\rightarrow 0$. At the same time, when this value of the volume-integrated tortuosity drop is not normalized by $T_{\Omega}(Re_{k}\rightarrow 0)-1$, it does not behave monotonically with porosity, i.e., the largest value of the drop is observed for the $\phi=0.59$ sample (see the inset of the top plot). Smaller values of the normalized $T_{\Omega}$ drop in the bottom subplot suggest the sooner dominance of the recirculation zones-related phenomena over the ones related to the percolating volume. The positions of the $T_{\Omega}$ minima increase with porosity, which is due to the reference length $\sqrt{k_{0}}$ in the Reynolds number changing by orders of magnitude between the samples (cf. Eq. (14)). When the diameter of the spherical obstacle is taken as the reference length (cf. Table 3), the profiles of $T_{\Omega}(Re_{k})$ fall more into a single curve, see Fig. 24. On the other hand, what the choice of different reference lengths in the Reynolds number cannot change is the range of $Re$ over which $T_{\Omega}$ is decreasing (in terms of decades, i.e., the width of the decreasing part of $T_{\Omega}(Re_{k})$ plot on logarithmic $Re_{k}$-axis). This range increases with increasing porosity. This suggests that as the porosity decreases and the geometrical confinement increases, the inertial effects appear in the system less abruptly.
Figure 14: Volume-integrated tortuosity for all SC samples. The top plot shows the raw values of $T_{\Omega}$ with the inset showing the same data but in different values range. The bottom plot shows the values rescaled by $T_{\Omega}$ at vanishingly small $Re_{k}$.
@@
We next look at the Reynolds number dependence of $T_{\Omega}$ and $T_{s}$ for SC samples, shown in Fig. 15. In the Darcy regime ($Re_{k}\rightarrow 0$), the values of both tortuosities stay constant and are compliant with each other, as shown in [27]. We attribute small visible discrepancies between the two tortuosities to the errors introduced during the generation of streamlines and the identification of percolating/recirculation volumes. When the Reynolds number increases, the streamline-based tortuosity, $T_{s}$, generally decreases for all systems, which was observed in stochastic and periodic porous media by other authors [22, 15]. The exception from this behavior is visible in the last decade of $Re_{k}$ in the case of $\phi=0.59$ sample, however even in this case the eventual increase of $T_{s}$ is small compared to the increase of $T_{\Omega}$ in the same range of Reynolds number. On the other hand, when the volume-based tortuosity, $T_{\Omega}$, is concerned, the inertial effects manifest themselves differently depending on the porosity, similarly to the stochastic porous media (cf. Fig. 7). The overall tendency is that $T_{\Omega}$ starts to deviate significantly from $T_{s}$ earlier when the porosity is decreased. For instance, in the lowest-porosity case, its value decreases slightly in the range $Re_{k}\in[3\cdot 10^{-3};10^{-2}]$ to later increase significantly until the steady-state regime ends. Such behavior resembles the one observed for the lower-porosity stochastic porous media in Fig. 7 and also for a real-life porous sample [16]. At the same time, for the highest-porosity sample, the volume-based tortuosity follows the streamline-based one closely and diverges from it only at $Re_{k}\approx 10^{3}$. Similarly to the system, the initial decline of $T_{s}$ is followed by its rise, however to a much smaller extent, . This is in line with the behavior of $T_{\Omega}$ for the higher-porosity stochastic porous media shown in Fig. 7, and in [14].
@@
The decreasing behavior of the streamline-based tortuosity can be attributed to the straightening of the percolating streamlines, which follows the growth of the recirculation zones volumes in the cavities and obstacles wakes. The rise of $T_{s}$ for $Re_{k}>2$ in the case of $\phi=0.59$ sample indicates that perhaps a non-trivial reorganization of the percolating streamlines shapes takes place there, reminiscent of those observed in porous media of more complex geometries [16, 44]. The volume-integrated tortuosity exhibits a more complex behavior, dependent on the sample’s porosity. Eq. (13) shows that the difference between $T_{\Omega}$ and $T_{s}$ is related to the integral of velocity magnitude inside the recirculation zones. Thus, we will further investigate the fall and the rise of the volume-based tortuosity in terms of the volume of the recirculation zones and the field of the velocity magnitude therein.
@@
Figure 15: $T_{\Omega}$ (circles) and $T_{s}$ (squares) for SC samples. The inset in the $\phi=0.1$ plot shows the magnification of the data to the range of $T_{s}$ values. The horizontal dashed lines represent the values of $T_{\Omega}$ calculated for vanishingly small $Re_{k}$.
@@
III.3 Origin of the decrease and rise of the volume-integrated tortuoisty – volume of the recirculation zones
@@
Let us focus on the numerator of the recirculation volume-related term from Eq. (13), $I[V]_{v}/I[V_{0}]_{p}$. For convenience, we divide it by the total fluid volume, $I[1]$, i.e.
@@
where the first term on the right-hand side represents the vortex volume normalized by the total volume of fluid, and the other one – the total velocity magnitude per unit recirculation zone volume. We consider each of those two components separately. Figs. 1619 show the visualization of the percolating volumes or recirculation zones for SC samples, for two different Reynolds numbers each. In all systems, as $Re_{k}$ rises, the recirculation zones grow in volume. In the case of $\phi=0.1$ system (Fig. 16), it is visible through the percolating volume shrinking to a straight jet connecting the inlet and the outlet. $\phi=0.59$ system behaves similarly (Fig. 17), although because there is more space between the obstacles in the neighboring periodic cells, the high-$Re_{k}$ streamlines still feature a slight curvature (cf. Fig. 11). In the two higher-porosity systems, there is no recirculation in the obstacles’ wakes for low $Re_{k}$, but with the increasing velocity they emerge, and in the case of $\phi=0.93$ sample (Fig. 18) they fill the whole fluid volume between the spheres in the streamwise direction at higher $Re_{k}$. We note that for low $Re_{k}$ in $\phi=0.26$ sample, a thin recirculation zone around the $y=z=0.5$ line is visible. However, the later examination of the velocity fields (cf. Fig. 12) confirmed that the recirculation zone is not present there, and its appearance in the histograms is most likely due to the numerical errors introduced in the streamlines generation.
To quantify the changes in the size of the recirculation zones, we plot their volume, $I[1]_{v}$, normalized by the total volume of fluid, $I[1]$, in Fig. 20. In $\phi=0.1$ system, there already exist recirculation zones in the Darcy regime (in the channels perpendicular to the main flow, cf. top visualization in Fig. 10). The increase in their volume starts at $Re_{k}\approx 2\cdot 10^{-3}$, which corresponds to the start of $T_{\Omega}$ decrease in Fig. 15. The vortex volume then continues to grow until the end of the steady state regime, with significant slowdown for $Re_{k}>4\cdot 10^{-2}$, the value at which the slope of the $T_{\Omega}$ rise in Fig. 15 decreases. This reduction in the recirculation volume growth pace indicates the strong geometrical confinement for the vortex development. We highlight that at the highest $Re_{k}$, there already exist secondary vortices in the perpendicular channels (cf. bottom visualization in Fig. 10). The behavior of the vortex volume and $T_{\Omega}$ in the $\phi=0.59$ sample is similar, with the start of rise of the vortices volume and fall of the tortuosity at $Re_{k}=10^{-1}$, and a visible decrease of the slope in both quantities visible for the highest-$Re_{k}$ case.
@@
On the other hand, in the two higher-porosity systems, there is no recirculation in the Darcy regime. Contrary to the two lower-porosity samples, there is no visible decrease in the slope of the recirculation zone volume growth in the highest Reynolds numbers. This might be caused by a lower geometrical confinement of the recirculation volumes when the porosity is high. For $\phi=0.93$ sample, the growth of the recirculation zones volume starts at about $Re_{k}=2$, slightly after the start of $T_{\Omega}$ decrease. We note once again that for this sample, the non-zero value of the vortex volume for $Re_{k}\rightarrow 0$ in Fig. 20 is caused by numerical errors. For the highest-porosity sample, the increase of the recirculation volume size starts at $Re_{k}\approx 2\cdot 10^{2}$ (cf. top visualization in Fig. 13), which is about two decades later than the visible start of $T_{\Omega}$ decrease for this system in Fig. 15. This discrepancy can be related to the ‘formation of inertial cores’ [45] when the inertia already starts to change the velocity field, although no separation yet occurs. As the recirculation volume starts to grow, it does so with an increasing slope until the end of the steady-state regime (cf. bottom visualization in Fig. 13). The growth of the recirculation zones in the case of the highest-porosity SC systems was highlighted by other authors [19] as the reason for lower drag compared to the flow over a single obstacle.
Table. 1 shows the difference between the volume of the recirculation zone in the highest and lowest investigated $Re_{k}$ for each SC sample. While the difference relative to the fluid volume decreases with increasing porosity, it reaches a maximum in the intermediate porosity when calculated relative to the volume of the obstacle.
Finally, we note that as the considered SC systems have a regular and simple geometry, phenomena like the non-strictly-monotonic growth of the volume of the recirculation zones [44] related to the instantaneous reorganization of the percolating flow paths are not visible here. Nevertheless, the start of the growth of the recirculation zones volume can still be related to the deviation from Darcy’s law for the lowest-porosity system (compare the data for $\phi=0.1$ in Figs. 20 and 22). Interestingly, no oscillations can be seen as well in the values of $\rho^{-}$ [16, 14] – the fraction of fluid volume with negative $V_{0}$ – in the case of the stochastic samples (results not shown, but are available in the repository, see Sec. VII), which suggests that their geometry might still not be complex enough for the spontaneous reorganization of the percolating flow paths to occur.
@@
Figure 16: Visualizations of the percolating volumes in $\phi=0.1$ SC system for the same values of $Re_{k}$ as in Fig. 10. The gray surfaces are the walls of the obstacles, and only a few chosen walls are shown for clarity. The visualizations use the histograms from which the percolating volumes were found. .
@@
Figure 17: Visualizations of the recirculating volumes for $\phi=0.59$ SC system for the same values of $Re_{k}$ as in Fig. 11. The gray surfaces are the walls of the obstacles. The visualizations use the histograms from which the recirculation volumes were found.
Figure 18: Visualizations of the recirculating volumes for $\phi=0.93$ SC system for the same values of $Re_{k}$ as in Fig. 12. The gray surfaces are the walls of the obstacles. The visualizations use the histograms from which the recirculation volumes were found.
Figure 19: Visualizations of the recirculating volumes for $\phi=0.999$ SC system for the same values of $Re_{k}$ as in Fig. 13. The gray surfaces are the walls of the obstacles. The visualizations use the histograms from which the recirculation volumes were found. .
Figure 20: Relative vortex volumes, $I[1]_{v}/I[1]$, for the studied SC systems.
$I[1]_{v}(Re_{k,\text{max}})-I[1]_{v}(Re_{k}\rightarrow 0)$
$\phi$
Rel. to fluid
volume $\%$]
Rel. to obstacle’s
volume $\%$]
$0.1$ 56 6.2
$0.59$ 20 28
$0.93$ 8.8 117
$0.999$ 0.07 63
Table 1: The difference between the volume of the recirculation zone at the minimal and maximal investigated Reynolds number for SC samples. Data are given relative to the volume of fluid and the solid obstacle.
@@
III.4 Origin of the decrease and rise of the volume-integrated tortuoisty – energy confinement in the recirculation zones
The increase of the recirculation zone volume can explain the decrease of $T_{\Omega}$ because it directly implies the straightening of all the percolating streamlines and thus the decrease of $T_{s}$ in Eq. (13). However, it fails to address the rise of the volume-based tortuosity at the highest Reynolds numbers. This is why now we turn our attention to the the other term in Eq. (15), the integral of the velocity magnitude inside the vortex per vortex volume, $I[V]_{v}/I[1]_{v}$ (further in the text, we shall refer to this quantity as the specific velocity magnitude in vortex, for brevity). Its dependence on the Reynolds number is shown in Fig. 21 for all SC systems. We note that since there is no recirculation zone found in the highest-porosity SC system in the Darcy regime, we assume that the value of the specific velocity magnitude in the vortex is zero in those cases.
@@
For all SC systems, the specific velocity magnitude in vortices increases with Reynolds number. In the two lower-porosity samples, the rise proceeds from the lowest $Re_{k}$ with the first-order slope and experiences a jump in the slope around $Re_{k}=10^{-2}$ for $\phi=0.1$ and $Re_{k}=4\cdot 10^{-1}$ for $\phi=0.59$. It later attains the slope of the order approximately one again, in the range of $Re_{k}$ corresponding to those where the growth of $T_{\Omega}$ slows down (cf. Fig. 15). The initial first-order dependence is explained by the linearity of the Stokes equation and proportionality between the body force magnitude $g$ and the mean flow velocity $\langle\boldsymbol{V}\rangle$ in the Darcy regime (i.e., when $g$ is increased, the velocity field inside the vortex is only scaled proportionally by a constant). The subsequently higher slope indicates that a non-trivial mechanism involving the momentum transfer to the recirculation zone and the growth of its volume comes into play in the inertial regime. At the same time, an analogy between the Darcy regime and the part of the inertial regime where the specific velocity magnitude in vortex decreases its slope again can be drawn in the sense that both are characterized by the recirculation zone growing more in kinetic energy than volume. This further highlights the importance of the geometrical confinement in the emergence of inertial effects in low-porosity samples [16].
In the two higher-porosity systems, the start of the rapid growth of the specific velocity magnitude in the vortex happens roughly at $Re_{k}$ when the recirculation zone volume starts to grow (cf. Fig. 20). It is interesting that even though the integral of the velocity magnitude over the whole fluid volume, $I[V]$, achieved at the highest Reynolds number is about $40$ times larger in the highest-porosity system ( 1855 vs. 47, see Table 2), the specific velocity magnitude in vortex is about 2 times higher in the lowest-porosity system ($\approx 140$ vs. $\approx 70$, see also Table 2). This suggests that strong geometrical confinement leads to the accumulation of relatively large amounts of the kinetic energy in the recirculation zones, quantified also by other authors using the variance of the kinetic energy [13, 15, 14, 16].
For maximal $Re_{k}$
$\phi$ $I[V]$ $I[V]_{v}/I[1]_{v}$
$0.1$ 47 142
$0.59$ 138 14
$0.93$ 298 27
$0.999$ 1855 70
Table 2: The integral of the velocity magnitude in the whole fluid, $I[V]$ and the specific velocity magnitude in vortex, $I[V]_{v}/I[1]_{v}$, for the studied SC systems at the highest Reynolds number investigated for each porosity.
In the end, at high Reynolds numbers, as the porosity of either stochastic or SC system decreases, more kinetic energy stays confined in the recirculation zones and the recirculation zones occupy a larger fraction of the fluid volume. Those two phenomena, reflected in the growth of the terms $I[V]_{v}/I[1]_{v}$ and $I[1]_{v}/I[1]$, respectively, lead to a greater increase of $T_{\Omega}$ close to the steady-state regime limit, compared to the higher-porosity samples. The dynamics of the growth of the recirculation zones also determine the amount of the fall of $T_{\Omega}$ between its $Re_{k}\rightarrow 0$ value and its minimum. A more pronounced growth of the volume and kinetic energy inside them in the lower-porosity samples leads to a sooner domination of the $I[V]_{v}/I[V_{0}]_{p}$ term over the $T_{s}$ term in Eq. (13). This, in turn, causes the divergence between the two tortuosities to begin at their values closer to $T_{\Omega}(Re_{k}\rightarrow 0)\equiv T_{s}(Re_{k}\rightarrow 0)$, rather than the minimum of $T_{s}$, achieved at high Reynolds numbers, which explains the tendencies visible in the bottom subplot of Fig. 14 and in Fig. 15.
@@
Figure 21: Integral of the absolute value of velocity in the vortex per vortex volume for SC systems (for brevity called the specific velocity magnitude in vortex in the text, $I[V]_{v}/I[1]_{v}$). The insets in the two topmost subplots show the same data as the main subplot but on a log-log scale.
IV Conclusions